What is the value of a preferred stock when the dividend rate is 16 percent on a $75 par value?


Problem 8-1

(Preferred stock valuation) What is the value of a preferred stock when the dividend rate is 16 percent on a $75 par value? The appropriate discount rate for a stock of this risk level is 14 percent. 

The value of the preferred stock is _______ . (Round to the nearest cent.)


(Preferred stock valuation) The preferred stock of Gandt Corporation pays a $0.50 dividend. What is the value of the stock if your required return is 11 percent? 

The value of the preferred stock is $_______  per share. (Round to the nearest cent.)


(Preferred stock valuation) Haney, Inc.'s preferred stock is selling for $23.25 per share in the market and pays a $2.25 annual dividend. 

a. What is the expected rate of return on the stock? 

b. If an investor's required rate of return is 12 percent, what is the value of the stock for that investor? 

c. Should the investor acquire the stock? a. The expected rate of return on the stock is % (Round to two decimal places.)



Problem 8-6 

(Preferred stock valuation) You are considering an investment in one of two preferred stocks, TCF Capital or TAYC Capital Trust. TCF Capital pays an annual dividend of $4.72, while TAYC Capital pays an annual dividend of 4.95. If your required return is 12 percent, what value would you assign to the stocks? 

The value of the TCF Capital preferred stock is $_______ per share. (Round to the nearest cent.)



Problem 8-8

(Common stock valuation) Mosser Corporation common stock paid $2.33 in dividends last year and is expected to grow indefinitely at an annual 3 percent rate. What is the value of the stock if you require a return of 11 percent? 

The value of the Mosser Corporation common stock is $ _______ (Round to the nearest cent)
Answer #1

Part 8.1:

Value of Preferred Stock = Preference Div / Required Ret

= [ Par Value * Div Rate ] / Required Ret

= [ $ 75 * 16% ] / 14%

= $ 12 / 14%

= $ 85.71

8.1 . 2:

Value of Preferred Stock = Preference Div / Required Ret

= [ $ 0.5 ] / 11%

= $ 0.5 / 11%

= $ 4.55

8.1.3:

8.1.3.a

Expected Ret = Div / Selling Price

= $ 2.25 / 23.25

= 0.0968 i.e 9.68%

8.1.3.b:

Value of Stock = Div / Required Ret

= $ 2.25 / 12%

= $ 18.75

8.1.3.c:

Stock is not adviced to buy as Selling price is 23.25 and Fair price is 18.75

8.6

Price of TCF Capital = Div / Required Ret

= 4.72 / 12%

= 39.33

Price of TAYC Capital = Div / Required Ret

= 4.95 / 12%

= 41.25

8.8:

Value of Common stock = D0(1+g)/ [ ke - g ]

= 2.33 ( 1 + 0.03) / [ 11% - 3% ]

= 2.33*1.03 / 8%

= 2.4 / 8%

= $ 30.00

Pls comment, if any further asssistance is required.